OpenStudy (anonymous):

Consider a quadratic formula y=ax^(2)+80x+1000 where a is an integer. (a) If the minimum value of y is a positive integer, suggest three possible values of a. (b) If the maximum value of u is at least 1200, suggest three possible values of a.

OpenStudy (anonymous):

another question: \[\log_{2} x+\log_{x} 2=\frac{ 5 }{ 2 }\] find x.

OpenStudy (raden):

because it's a minimum function, so a > 0 and because the minimum value of y is a positive it must be a definite positive

OpenStudy (anonymous):

question 3: If \[9^{\frac{ 1 }{ h }} = 11^{\frac{ 1 }{ k }} =99\], where h and k are not equal to 0, determine the value of h+k without finding the values of h and k.

OpenStudy (anonymous):

I have already known a > 0 .....

OpenStudy (raden):

but D < 0

OpenStudy (anonymous):

D ??

OpenStudy (raden):

lol, D=discriminant = b^2-4ac

OpenStudy (anonymous):

.....and then?

OpenStudy (raden):

D < 0 b^2 - 4ac < 0 (80)^2 - 4a(1000) < 0 6400 - 4000a < 0 a > 1.6 so, the intersection between a > 0 and a > 1.6 is a > 1.6 because a an integer number, u can take a=2,3,4,5, e.t.c

OpenStudy (anonymous):

oh ya, I got it. how about part b? There are two parts...

OpenStudy (raden):

hmmm.... i dont understand why there is "u" there

OpenStudy (raden):

@algebraic, do u know what mean u there ??

OpenStudy (anonymous):

it should be y instead of u. sorry for that

OpenStudy (raden):

ohhhh... :P

OpenStudy (anonymous):

I have a question in part a , I substitute a=2, then the equation is : y=2x^(2)+80x+1000. How can I be factorize?

OpenStudy (raden):

well, if it's a maximum function, so a < 0 and at least 1200, it means y >= 1200 (with y = -D/4a) -(b^2 - 4ac)/4a > = 1200 (b^2 - 4ac)/4a <= 1200 ((80)^2 - 4a(1000))/4a <= 1200 (6400 - 4000a)/4a <= 1200 (1600 - 1000a)/a <= 1200 (16 - 10a)/a <= 12 for zeros, a=0 or a=1.6